Find the LCD for the given rational expressions and convert each rational expression into an equivalent rational expression with the LCD as the denominator.
5x/x^2-1, -4/x^2-2x+1

Answer by nyc_function(2741)

The LCD for the fraction 5x/(x^2-1) = (x - 1)(x + 1).
The LCD for the fraction -4/(x^2-2x+1) = (x - 1)(x - 1).